x^2 - 4 / x^2 + 12x +27 =
^=to the power of, I don't know how to use my keyboard to make the correct symbols. This is a division problem that I need help with.

Found 2 solutions by helpnalgebra, london maths tutor:
Answer by helpnalgebra(91)

x^(2)-(4)/(x^(2))+12x+27
x^(2)*(x^(2))/(x^(2))+12x*(x^(2))/(x^(2))-(4)/(x^(2))+27*(x^(2))/(x^(2))
(x^(4))/(x^(2))+(12x^(3))/(x^(2))-(4)/(x^(2))+(27x^(2))/(x^(2))
(x^(4)+12x^(3)-4+27x^(2))/(x^(2))

(x^(4)+12x^(3)+27x^(2)-4)/(x^(2)) answer

Answer by london maths tutor(243)

x^2 - 4 = (x+2)(x-2)
x^2 + 12x + 27 = (x+9)(x+3)
Answer: [(x+2)(x-2)]/[(x+9)(x+3)]